Beam Test Meeting Mon. Mar. 26, 2012

From GlueXWiki
Jump to: navigation, search


Hall B doing electron running through ~April 1

reactivation during the day, trying to go to 5 pass photon this evening

Integrated Hall B run successful, identified 3 energies

cable guide change under tagger in next few days


resolution at 15.6 % for ~220 MeV electrons

still waiting on 5 pass beam


start at 22.8 % point

align to counter using angle and straight edge

start with detector perpendicular, better understood

move to additional realistic angle later

worked through several dead channels

2 shorted cables

problem with u board

eventually everything was fixed

can control HV from epics

building full DAQ setup now